# Federated Unlearning Made Practical: Seamless Integration via Negated Pseudo-Gradients

### v1

@@ V-D Main Results
Setting Algorithm Rounds ($\downarrow$) CE ($\uparrow$) Test Acc. Forget Acc. MIA [38] MIA [39] Multiple Unlearning, CIFAR-100, Non-IID, $E$=1, ResNet-18 Original 53.81±0.0 63.06±6.83 75.95±7.55 61.09±8.32 Retrain 47.32±2.13 30.09±3.29 39.66±3.90 27.89±3.12 NoT [19] 6.6±1.82 30.0$\times$ 48.06±2.24 39.16 (9.07 ±2.83) 50.51 (11.26 ±3.39) 41.8 (13.9 ±4.52) PUF-Special 3.2±1.1 62.5$\times$ 48.53±2.14 34.12 (4.04 ±3.69) 46.09 (6.64 ±3.46) 33.5 (5.61 ±4.21) PUF-Regular 8.0±2.83 25.0$\times$ 47.85±2.3 31.24 (2.87 ±1.86) 42.85 (3.89 ±2.94) 31.35 (3.78 ±3.13) Multiple Unlearning, CIFAR-100, Non-IID, $E$=1, MiT-B0 Original 75.03±0.0 84.60±5.25 76.84±8.95 73.8±1.40 Retrain 70.86±1.32 54.30±3.49 45.36±3.64 43.5±1.61 NoT [19] 3.2±3.03 15.6$\times$ 71.38±0.94 69.89 (15.58 ±6.29) 62.52 (16.55 ±8.89) 58.95 (15.45 ±7.50) PUF-Special 6.0±1.22 8.3$\times$ 71.28±1.44 60.58 (6.28 ±2.08) 52.97 (7.25 ±2.71) 50.50 (7.01 ±1.84) PUF-Regular 16.0±3.04 3.1$\times$ 70.87 ±1.49 56.57 (2.26 ±1.07) 48.67 (2.55 ±2.32) 45.62 (2.10 ±1.12)
TABLE IV: Performance of PUF and other baselines for multiple unlearning (unlearning two clients simultaneously). The column meaning is the same as Table III. Lower $\Delta$ corresponds to better unlearning performances.
@@
Multiple Unlearning. Table IV presents the performance on federated CIFAR-100 (non-IID) using a ResNet-18 or an MiT-B0 architecture, where a pair of randomly chosen clients was designated as unlearning clients (averaged over five independent experiments). We do not include PGA as a baseline because it is not directly extensible to unlearning multiple tasks. This limitation arises from its reliance on a reference model, which is obtained by removing the last stored model update of the target client. As a result, the mutual dependency of unlearning updates prevents its straightforward adaptation to multiple unlearning requests.
As shown in Table IV, the most effective unlearning of multiple clients simultaneously is achieved by either PUF-Special or PUF-Regular. While NoT benefits from a relatively short recovery phase, it demonstrates significantly weaker forgetting performance, as indicated by larger gaps in Forget Accuracy and MIAs. Notably, even when handling multiple unlearning requests, PUF-Regular, like NoT, operates without requiring any modifications to the standard FedAvg protocol.
Algorithm Rounds ($\downarrow$) CE ($\uparrow$) Test Acc. Forget Acc. Forget Loss Original 86.54±0.0 88.94±4.58 0.08±0.05 Retrained 82.69±1.68 73.29±7.23 0.22±0.10 NoT [19] 58.6±22.74 8.5$\times$ 83.43 ±0.74 77.37 (3.91±2.42) 0.17 (0.06±0.02) PUF-Special 41.8±11.95 12.0×12.0$\times$12.0 × 83.06±0.37 76.71 (3.38±2.27) 0.17 (0.05±0.02)
TABLE V: Performance of PUF and other baselines on ProstateMRI dataset. Values in parenthesis represent the mean absolute difference with the retrained model, where a lower value is better.
Task Agnosticism. As highlighted in Sections II and IV, the design of PUF ensures that the unlearning phase remains task-agnostic, as both the unlearning and remaining clients follow the standard training procedure without any modifications to their local learning routines. Table V reports the results for the ProstateMRI experiments, which focus on a segmentation task rather than classification. PUF outperforms all baselines in terms of both recovery efficiency (i.e., shorter recovery phase) and forgetting efficacy. As discussed in Section V-E, PUF can be further tuned for improved forgetting performance by increasing the scaling factor $\eta_{u}$, albeit at the cost of a longer recovery phase. Notably, NoT does not offer this tuning flexibility.
@@
Figure 6: Performance of PUF with varying hyper-parameters ($\eta_{u}$, $E_{u}$) compared to related methods. When $E_{u}$ is omitted, it is set to 1. X-axis: number of recovery rounds required to match the retrained model’s test accuracy. Y-axis: gap in forget accuracy compared to the retrained model. Points closer to the origin indicate better performance. The experimental setting is indicated in subcaption as a triple dataset, data distribution, model architecture. The Natural baseline reports results for a naive strategy of fine tuning the global model without the participation of the unlearning client, no explicit unlearning is performed.
Figure 7: Performance of PUF with varying hyper-parameter $\eta_{u}$ compared to related methods for the ProstateMRI setting.
@@ V-E Hyper-parameter Tuning of PUF
Notably, increasing local epochs ($E_{u}>1$) for PUF reduces $\Delta$ Forget Accuracy but requires more recovery rounds compared to $E_{u}=1$ with the same $\eta_{u}$. Interestingly, increasing the scaling factor $\eta_{u}$ while keeping $E_{u}=1$ achieves similar improvements without additional local computational overhead for unlearning clients, making it more practical in resource-constrained settings. Both PUF-Regular and PUF-Special achieve high effectiveness by simply tuning the scaling factor ($\eta_{u}$) while maintaining $E_{u}=1$. As shown in Figure 6, PUF demonstrates low sensitivity to the $\eta_{u}$ hyper-parameter, resulting in a cluster of well-performing configurations.
In contrast, we observe that PGA requires tuning multiple hyper-parameters – local learning rate, local epochs, an early stopping threshold, and gradient-clipping threshold – while also assuming full client participation and stateful clients. On the other hand, NoT does not require specific hyperparameter tuning but fails to achieve significant forgetting. As shown in Figure 6.a, NoT performs only slightly better than the Natural baselines in both recovery efficiency and forgetting effectiveness. This highlights PUF’s trade-off between simplicity, ease of hyperparameter optimization, and unlearning performance compared to other baselines.
@@ VI Conclusive Remarks
In this paper, we introduced PUF, a novel FU method that leverages negated pseudo-gradients to enable clients to exercise their right to be forgotten. Aligned with the design principles of FL, PUF does not require storing any processed client information and relies only on ephemeral updates, even during unlearning. It integrates seamlessly with standard FedAvg, by requiring no significant modifications or additional computational/communication overhead. As a result, PUF is both task-agnostic and capable of handling multiple unlearning requests concurrently, which are two critical aspects that most existing literature does not consider. Experimental results on two classification datasets and a segmentation task, across varying degrees of data heterogeneity and different model architectures, are reported to show the general applicability of the proposed approach: PUF has clearly demonstrated to be more efficient in recovering expected performance and more accurate in inducing forgetting of requested data compared to state-of-the-art baselines.

### v2

@@ V-D Main Results
Setting Method Efficacy Efficiency Test Acc. Forget Acc. ($\Delta$ $\downarrow$) MIA[Song] ($\Delta$ $\downarrow$) MIA[Yeom] ($\Delta$ $\downarrow$) Communication Computation Storage Bytes ($\times$ $\uparrow$) FLOPs ($\times$ $\uparrow$) Bytes ($\times$ $\uparrow$) Multiple Unlearning, CIFAR-100, Non-IID, ResNet-18, EE=1 Original 53.8±0.0 63.1±6.8 76.0±7.6 61.1±8.3 Retrain 47.3±2.1 30.1±3.3 39.7±3.9 27.9±3.1 1.44e11 (1.0×) 1.20e15 (1.0×) 4.49e07 (1.0×) FedAU 48.3±2.3 35.6 (5.7±2.4) 45.6 (6.4±3.6) 33.6 (5.7±4.1) 7.90e09 (18.2×) 6.60e13 (18.2×) 4.49e07 (1.0×) NoT 48.1±2.2 39.2 (9.1±2.8) 50.5 (11.3±3.4) 41.8 (13.9±4.5) 4.74e09 (30.4×) 3.96e13 (30.3×) 4.49e07 (1.0×) PUF-Special 48.5±2.1 34.1 (4.0±3.7) 46.1 (6.6±3.5) 33.5 (5.6±4.2) 2.48e09 (58.1×) 2.07e13 (58.0×) 4.49e07 (1.0×) PUF-Regular 47.9±2.3 31.2 (2.9±1.9) 42.9 (3.9±2.9) 31.4 (3.8±3.1) 7.54e09 (19.1×) 5.55e13 (21.6×) 4.49e07 (1.0×) Multiple Unlearning, CIFAR-100, Non-IID, MiT-B0, EE=1 Original 75.0±0.0 84.6±5.3 76.8±9.0 73.8±1.4 Retrain 70.9±1.3 54.3±3.5 45.4±3.6 43.5±1.6 1.06e10 (1.0×) 3.40e15 (1.0×) 1.33e07 (1.0×) FedAU 71.4±0.9 58.9 (4.6±2.3) 50.1 (4.7±2.3) 47.2 (3.7±1.9) 3.19e09 (3.3×) 1.02e15 (3.3×) 1.33e07 (1.0×) NoT 71.4±0.9 69.9 (15.6±6.3) 62.5 (16.6±8.9) 59.0 (15.5±7.5) 6.80e08 (15.6×) 2.18e14 (15.6×) 1.33e07 (1.0×) PUF-Special 71.3±1.4 60.6 (6.3±2.1) 53.0 (7.3±2.7) 50.5 (7.0±1.8) 1.12e09 (9.5×) 3.57e14 (9.5×) 1.33e07 (1.0×) PUF-Regular 70.9±1.5 56.6 (2.3±1.1) 48.7 (2.6±2.3) 45.6 (2.1±1.1) 3.93e09 (2.7×) 1.17e15 (2.9×) 1.33e07 (1.0×)
TABLE V: Performance of PUF and other baselines across different settings for multiple unlearning when two clients request forgetting. For efficacy metrics, baseline results are expressed as mean metric value (mean $\Delta$ $\pm$ standard deviation), with $\Delta$ in parentheses representing the average absolute difference from the Retrain baseline. Lower $\Delta$ corresponds to better unlearning performances. For efficiency metrics, higher $\times$ (reduction over Retrain baseline) are better. The best-performing method is shown in bold.
[…]
Setting Method Efficacy Efficiency Test Acc. Forget Acc. ($\Delta$ $\downarrow$) Forget Loss ($\Delta$ $\downarrow$) Communication Computation Bytes ($\times$ $\uparrow$) FLOPs ($\times$ $\uparrow$) ProstateMRI, U-NET Original 86.5±0.0 88.9±4.6 0.08±0.05 Retrain 82.7±1.7 73.3±7.2 0.22±0.10 2.74e11 (1.0×) 5.49e16 (1.0×) PGA 82.9±0.4 77.0 (3.8±2.3) 0.18 (0.06±0.02) 2.69e10 (10.2×) 5.44e15 (10.1×) NoT 83.4±0.7 77.4 (3.9±2.4) 0.17 (0.06±0.02) 2.63e10 (10.4×) 5.27e15 (10.4×) PUF-Special 83.1±0.4 76.7 (3.4±2.3) 0.17 (0.05±0.02) 2.30e10 (11.9×) 4.62e15 (11.9×)
TABLE VI: Performance of PUF and other baselines across for ProstateMRI segmentation task. For efficacy metrics, Lower $\Delta$ corresponds to better unlearning performances. For efficiency metrics, higher $\times$ (reduction over Retrain baseline) are better. The best-performing method is shown in bold.
@@
Multiple Unlearning. Table V presents the performance on federated CIFAR-100 (non-IID) using a ResNet-18 or an MiT-B0 architecture, where a pair of randomly chosen clients was designated as target clients (averaged over five independent experiments). We do not include PGA as a baseline because it is not directly extensible to unlearning multiple tasks. This limitation arises from its reliance on a reference model, which is obtained by removing the last stored model update of the target client. As a result, the mutual dependency of unlearning updates prevents its straightforward adaptation to multiple unlearning requests. Similarly, we do not include MoDe, since its memory-guidance phase is not easily adaptable to this scenario.
[…]
@@
Figure 5: Performance of PUF with varying hyper-parameters ($\eta_{u}$, $E_{u}$). When $E_{u}$ is omitted, it is set to 1. X-axis: number of recovery rounds required to match the retrained model’s test accuracy. Y-axis: gap in forget accuracy compared to the retrained model. Points closer to the origin indicate better performance. The experimental setting is indicated in subcaption as a triple dataset, data distribution, model architecture. The Natural baseline reports results for a naive strategy of fine tuning the global model without the participation of the unlearning client, no explicit unlearning is performed.
Task Agnosticism. As highlighted in Sections II and IV, the design of PUF ensures that the unlearning phase remains task-agnostic, as both the unlearning and remaining clients follow the standard training procedure without any modifications to their local learning routines. Table VI reports the results for the ProstateMRI experiments, which focus on a segmentation task rather than classification. FedAU is not applicable to segmentation tasks like Res-UNet, as it requires a separate classification head, which such architectures lack. MoDe’s memory-guidance phase is also not clearly adaptable to non-classification tasks, with no instructions in the original paper. PUF outperforms all baselines in terms of both recovery efficiency (i.e., shorter recovery phase) and forgetting efficacy. As discussed in Section V-E, PUF can be further tuned for improved forgetting performance by increasing the scaling factor $\eta_{u}$, albeit at the cost of a longer recovery phase. Notably, NoT does not offer this tuning flexibility.
@@ V-E Hyper-parameter Tuning of PUF
Notably, increasing local epochs ($E_{u}>1$) for PUF reduces $\Delta$ Forget Accuracy but requires more recovery rounds compared to $E_{u}=1$ with the same $\eta_{u}$. Interestingly, increasing the scaling factor $\eta_{u}$ while keeping $E_{u}=1$ achieves similar improvements without additional local computational overhead for target clients, making it more practical in resource-constrained settings. Both PUF-Regular and PUF-Special achieve high effectiveness by simply tuning the scaling factor ($\eta_{u}$) while maintaining $E_{u}=1$. As shown in Figure 5, PUF demonstrates low sensitivity to the $\eta_{u}$ hyper-parameter, resulting in a cluster of well-performing configurations.
In contrast, we observe that other baseline methods require multiple hyper-parameter to tune. For example, PGA necessitates adjusting the local learning rate, local epochs, an early stopping threshold, and gradient-clipping threshold, while also assuming full client participation and stateful clients. MoDE involves tuning the number of memory-guidance and degradation rounds, separate learning rates for each, and a degradation parameter controlling momentum. FedAU needs to tune the weight to assign to the auxiliary module when merging it with the original model, as well as the number of epochs to train the module. On the other hand, NoT does not require specific hyperparameter tuning but fails to achieve significant forgetting, as shown in Table III. This highlights PUF’s trade-off between ease of integration, ease of hyperparameter optimization, and unlearning performance compared to other baselines.
Figure 6: Performance of PUF with varying hyper-parameter $\eta_{u}$ for the ProstateMRI setting.
@@ VI Conclusion
In this paper, we introduced PUF, a novel FU method that leverages negated pseudo-gradients to enable clients to exercise their right to be forgotten. Aligned with the design principles of FL, PUF does not require storing any processed client information and relies only on ephemeral updates, even during unlearning. It integrates seamlessly with standard FedAvg, by requiring no significant modifications or additional computational/communication overhead. As a result, PUF is both task-agnostic and capable of handling multiple unlearning requests concurrently, which are two critical aspects that most existing literature does not consider. Experimental results on two classification datasets and a segmentation task, across varying degrees of data heterogeneity and different model architectures, are reported to show the general applicability of the proposed approach. PUF has clearly demonstrated to be more efficient in recovering expected performance and more accurate in inducing forgetting of requested data compared to state-of-the-art baselines.
Appendix A Hyper-parameter Tuning
We compared PUF against five state-of-the-art baseline methods: FedEraser [16], PGA [19], FedAU [21], MoDe [20], and NoT [22]. Below, we summarize the hyper-parameter tuning for each method.
FedEraser [16]: we set the retention interval of rounds between stored updates to 1 (tuned in $\{1,2\}$) and the calibration epochs (number of local epochs during calibration training) to 0.5.
PGA [19]: we adopted the same hyper-parameters prescribed in the original paper, i.e., gradient clipping set to $5.0$, batch size in $\{512,1024\}$, local unlearning epochs set to $5$, and projection on the $\ell_{2}$-ball of radius $\delta$ around $w_{\text{ref}}$. We additionally tuned their early stopping threshold in the range $[2.2,10.0]$, selecting $9.0$ for the IID setting and $6.5$ for the Non-IID setting.
MoDe [20]: we set the number of memory-guidance rounds to $10$ and degradation rounds to $6$, the learning rate for memory guidance to $0.0005$, and the momentum parameter to $0.95$ (tuned in the range $[0.9,0.99]$).
FedAU [21]: we used $10$ local epochs to train the auxiliary module, and tuned the weight for auxiliary module integration in $[0.0,1.0]$, selecting $0.04$ as the best configuration.
NoT [22]: we negated the weights of the first layer of the neural network, following the prescription in the original paper.
PUF: for PUF-Regular, we set the scaling factor for the aggregated update from unlearning clients to $\eta_{u}=20.0$ and the learning rate for retained clients to $\eta_{r}=1.0$. For PUF-Special, we set $\eta_{u}=2.0$. Unless otherwise specified, we used SGD as the server-side optimizer with $\eta_{s}=1.0$.
Appendix B Cost Estimation Methodology
We report three complementary efficiency metrics: communication cost (in bytes), computational cost (in FLOPs), and storage overhead (in bytes). These metrics, preferred over wall‐clock time as in [22], are hardware- and implementation-agnostic, enabling fair comparisons across tasks, datasets, and systems. Unlike time, they isolate algorithmic efficiency from factors such as hardware heterogeneity or parallelization. We also report the relative improvement over the Retrain baseline. For each method, we account for both the unlearning phase and the recovery phase, with the latter corresponding to standard FedAvg rounds among $C_{r}$ clients.
Symbol Description
$B$ Bytes per parameter (4 for float32)
$P$ Model parameters
$P_{c}$ Classifier params (FedAU)
$C$ Total clients
$C_{u}$ Unlearning clients
$C_{r}$ Remaining clients $C-C_{u}$
$F$ FLOPs per image (0.150.15G ResNet-18, 1.71.7G MiT-B0)
NN Samples per client
$E$ Local epochs (generic)
$R_{\mathrm{ret}}$ Retained/calibration rounds (FedEraser)
$E_{\mathrm{cal}}$ Epochs per calibration round (FedEraser)
$E_{\mathrm{asc}}$ Epochs of ascent (PGA)
$R_{d},R_{m}$ Degradation and memory-guidance rounds (MoDe)
$R_{\mathrm{rec}}$ Recovery rounds
TABLE VII: Notation used for cost estimation formulas.
B-A Communication Cost
For each round, we compute the total bytes exchanged across participating clients (uplink + downlink) in both the unlearning phase and the recovery phase. The overall communication cost is obtained by summing the costs of these two phases. The general per‐round communication cost is:
$\mathrm{Comm}_{t}=2\,P\,B\,|\mathcal{S}_{t}|$
where $|\mathcal{S}_{t}|$ is the number of clients participating in round $t$.
For the recovery phase (same for all methods), the communication cost is:
$\mathrm{Comm}_{\mathrm{rec}}=2\,P\,B\,C_{r}\,R_{\mathrm{rec}}$
For the unlearning phase, the communication costs are method-specific, as reported in Table VIII.
B-B Computational Cost
The computational cost (in FLOPs) includes both the unlearning phase and the recovery phase. We generically estimate FLOPs in a training round as:
$\mathrm{FLOPs}=F\,N\,E\,C$
with $F$ denoting FLOPs per image.
For the recovery phase, the computational cost is:
$\mathrm{Comp}_{\mathrm{rec}}=F\,N\,E\,C_{r}\,R_{\mathrm{rec}}$
For the unlearning phase, the computational costs per method are shown in Table VIII.
B-C Storage Overhead
We report the maximum persistent storage (in bytes) required across clients and/or the server over multiple rounds and necessary for unlearning. The baseline cost for storing one global model is:
$\mathrm{Storage}_{\mathrm{baseline}}=P\,B$
which corresponds to the Retrain baseline following FedAvg. Storage requirements per method are summarized in Table VIII.
Method Communication Cost Computation Cost Storage Overhead FedEraser $2\,P\,B\,C_{r}\,R_{\mathrm{ret}}$ $F\,N\,E_{\mathrm{cal}}\,C_{r}\,R_{\mathrm{ret}}$ $C\,R_{\mathrm{ret}}\,P\,B$ PGA 2PB2\,P\,B $F\,N\,E_{\mathrm{asc}}\,C_{u}$ CPBC\,P\,B FedAU $2\,P_{c}\,B$ Negligible PBP\,B MoDe $2\,P\,B\,\big(C_{r}\,R_{d}+C\,R_{m}\big)$ $F\,N\,E\,\big(C_{r}\,R_{d}+C\,R_{m}\big)$ 2PB2\,P\,B PUF-Special $2\,P\,B\,C_{u}$ $F\,N\,E\,C_{u}$ PBP\,B PUF-Regular 2PBC2\,P\,B\,C FNECF\,N\,E\,C PBP\,B NoT Negligible Negligible PBP\,B
TABLE VIII: Communication, computation, and storage cost formulas for each unlearning method. Notation described in Table VII.